OpenStudy (anonymous):

determine the vertical asymptotes and the end behavior of the following function. 2x^2+3x+1/((x^2-2)(5-x^2))

OpenStudy (amistre64):

i dont see it canceling... got 2 linears up top and 2 irreduciles on the bottom

OpenStudy (amistre64):

x = +- sqrt(2) and +-sqrt(5) are your vas then

OpenStudy (amistre64):

ha is y=0

OpenStudy (amistre64):

its like i was never wrong ;)

OpenStudy (anonymous):

haha how did you get y=0? is it from the denominator ?

OpenStudy (amistre64):

it from the shortcut to finding has.... if the bottom degree is greater than the top degree; at all goes to 0 in the end

OpenStudy (anonymous):

so you expand the denomiator out first. right youll get 2X^2/x^4

OpenStudy (amistre64):

pretty much; the coeffs are meangingless in this since the x^4 overpowers everything else

OpenStudy (anonymous):

cool that was easy.. why does it look soo hard. LOL

OpenStudy (anonymous):

they make it seem that way to annoy you. big denominator means small number (by which i mean close to zero). think about \[\frac{(10^6)^2}{(10^6)^4}=\frac{10^{12}}{10^{24}}=\frac{1}{10^{12}}\] a number very close to zero